Tech Staff Meeting Notes September 5, 2006

Current Top Priorities are:LNXT Node BB upgrade, Station Maintenance, and ET Stations. (Also wish to get INV populated.)

There are four levels of priority.
PrioritiesWish to try to get these done (or make progress on) before the next meeting (yeah right)
Second TierFit this in too if an opportunity arises. May be good to do instead of priorities but check with Mitch. Likely to be a priority in a previous and/or subsequent month
As Time PermitsFit these in when you can but not instead of first or second priorities
OtherStuff we hope to get to and don't want to forget about

This redistributed task list agreed to at the November meeting. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything. Mitch is going to update this shortly.

Greg's September ToDo List is incorporated into this page.


Lenox Agreement. Jim noticed the 12 year agreement for the LNXT node is about to expire. Mitch will renew.

BB network upgrade.
Final two sensors are in the vault. See Mitch's PDF webpage for results.
MKTA and NMAD nodes are complete. Have LNXT node to go. HALT is going to require a taller freestanding tower at the station.

Notes from James:

MACC/CJTR - Maintenance on these has been done. We're working on getting a wireless shot off the base to some local school or business at CJTR. I don't expect any major developments here until next year.

PIGT/CVMS/MAIR - We'll soon have three new NetRS GPS receivers and will do a receiver swap for these three sites. This will take a maximum of two days and John-Paul or I can do this ourselves.

MAIR - The wireless Internet lead went bust. Apparently, the supplier in East Prairie isn't interested unless we cough up $750 and let _them_ buy the radio hardware. Safe to say that's _not_ going to happen. One-time setup fees are fine, but Bob's not interested in recurring charges for services.

RLAP - The air cells continue to hold but will need to be removed or replaced in the future. We'll use them until they die. The permission situation for the terminal building may have changed. I'd like RLAP to remain solar until we get the go-ahead to move our stuff back inside. We may need to use a second solar panel and battery for this as daylight hours fade.

STLE - Our plans for comms at the airport in Steele went south. It looks like we'll be using radios for a wireless shot between STLE and the school as originally planned. I'll be planning a day-trip to set up a radio test between the school and STLE.

MAIR/STLE/RLAP/PTGV - While these sites are still offline, we still have to use sneakernet to get their data. This has become part of my regular routine early each month, though I'd like to get John-Paul or Bill to help out, eventually. We'll download these sites locally until they get comms.

PTGV - The 5VDC power supply handling our router needs to be beefed up. I'll just buy one and do this work myself next time I'm out there.

CERI - We need to get an antenna setup here at CERI for GPS research and testing. Greg and I have determined that the roof of the FDL would be best, with an antenna cable run through a gable on House 3 and into a section of Paul Bodin's old space. This will wait for cooler weather.

John Paul has a project to install two new near-field stations at reelfoot. We've got the go-ahead for a site in New Markham. The second site should be north or west of Tiptonville, but not too far away. Ideally, RLAP, NWCC, and these two new sites should fall roughly along a line across the Reelfoot Scarp.

Radios - I've borrowed a pair of Freewaves from Jim and will be making radio tests at Tiptonville, Steele, and maybe East Prairie. This work should start ASAP. We need to determine if we can make wireless shots between our stations in Tiptonville, and also if the wireless shot over Steele will work. Fridays will be my field days this semester. I'll need to borrow Bill or Chris for this work, so I'll be sure to notify ahead of time.

Enclosures - Bob ordered enough hardware to build two more GPS enclosures. We'll need them assembled for New Markham and the other new site when the stuff arrives. I'll get with Greg about this.

Maintenance - Bob's negotiating with the UNAVCO Facility to have them download and QC our GPS data. Maintenance is not included in this package, so we'd still have to handle that. We'll be using HCES in Troy as the test site for this. Based upon the results, Bob will either decide to go ahead and turn other sites over, or to tell the Facility to get stuffed.


The map shows stations with yellow triangles being new ones.
SMNC may be getting interference and the gain channel is showing up on the vertical (but doesn't look right). John will visit this week, swap the radios, and inject a known signal on the pins.
SMNC4done (see note above)
BHT8move (eventually)
ETT is repaired. Has some breaks in the coax. Should probably replace entire length.
Flat gap, Van hill, and LRVA are on deck.
Need vehicle registration and insurance documents. Either the originals or something that says they are on file at State. Not a big deal in TN but sure is in other states. John as this odd aversion to jail cells. Chris is investigating.
TZTN needs a gate and have episensors swapped. The one there now belongs to SLU.
TRYN keeps blowing a 1 amp fuse.
GFM has a cable problem. ANSS.
8g to 4g is complete including and we are square with DTA.
Bill is getting familiarized with these so he can take them over.
Continued problems with mss100's losing connectivity (SMOC: simple matter of connectivity). The good news is Jim verified that the stations are still working when SMOC is lost.
John upgraded gps firmware at 2 of 3 gps. Mitch doesn't recall which ones. John says Knoxville and 2 in charleston still need upgraded. This is the list from before:
SWET-3H52, ADSC-5657, and CSTN-5637 need to have the leap second corrected. SHTN-5648 and TRSC-5662 might also but need a site visit to verify.
Mitch needs to contact UKY tech to get a site visit to MVKY.
ATTN is having firewall problems. Internet is apparently routed through somewhere in Nashville which is being reticent to allow CERI access. John will get contact info to Mitch.
MSS power cycle at ADSC didn't work. John will need to visit.
CVTN and ASAR optical isolation. Need to upgrade to year 3 hardware. Hardware ready. James brought back some nice pictures from CVTN to help figure out how to do it (e.g. the conduit is way too small).

Routine Maintenance.

Unscheduled Maintenance
Jim and Bill fixed a bunch of these. Mitch isn't sure which. Some of work fine in the shop; Greg will have a look. Here's Holly's list from the last meeting:
LNXT Node:
NMAD Node:
MKTA Node:

Second Tier Priorities

Need to follow up on ATV training.

NCCU Gordana Vlahovic is getting a cmg6td to install at North Carolina Central University. Invoice paid. Still owe here a vault (John has it) and installation advice. Need to get a public display. Scream to Scream from the datalogger pc to the display pc would be best. Don't forget to NOT allow remote config on display scream. John will be doing this later this month.

Steiner Brain Dump Seminar Gregs list is starting to take shape. While anyone is invited its geared to tech staff. Schedule will vary and be weather dependent to avoid field work interruptions. Will start within a few weeks.
How do we record this for posterity? Video would be enormous in size (potential of 104 hours or more of video). Audio might be useful but would not make a very good reference. Mitch thinks written is best preferably in html or pdf. Perhaps have someone (who) take everyones notes and combine them with Greg's notes and make a web page. Audio/video cache to help clarify might be useful for whoever does this.

FDL. Mitch needs to talk with Horton about what to do about FDL laptops. Clearly we're not getting help from Bodin.
Batteries are a problem. They they show up and dissappear, some get charged too long and others not at all. Bill will be the Battery Czar. Anyone using batteries needs to go through him. Bill will also put some labels up to help know who's are who's, what is charged, and what needs charging.

3904 Power Need to draft an updated wiring diagram of 3904. Need to update the diagrams on the circuit panels. Mitch will contact P-plant and ask to have someone at CERI involved when they do work to minimize disruption of mission critical systems.

As Time Permits

Mitch needs to write an ew module to monitor the environmental channels newly installed at the nodes. Done. Its a shell script that runs twice a day. Computes several minute averages. Emails when there's a problem. Its is /gaia/home/mwithers/Projects/ENVMonitoring.

Need to come up with a good way for John to view real data (not helicorders) on is pc over dsl. Mitch will send John info on ssh and exceed.

Generators Old generator needs an exhaust pipe.

Inventory Sid has Ketchum's INV program installed on smeagol. Jim got initial training. Need to show Bill how to use it. Jim will be the lead and Bill will do initial update. Routine updates can be done by whoever does the work (requires usernames and passwords in mysql). Not sure yet how best to handle dataless seed production. Mitch needs to learn the admin tasks. Startup script is /gaia/smeagol/local/bin/runINV.


Many of the air cells in storage are ancient and bad. Bill will go through them and dispose of the old. Greg has method for safe and legal disposal.

I-40 Bridge.
Boreholes are drilled and more or less secured. Need to do prep for new instrumention at HDBT and take care of drainage now that a big pile of stone and dirt was piled around the station. Greg's theory is lightning is clobbering gps. Has suppression constructed and ready to install.
Need a better pc at AZTN now that we have equivalent of two nodes plus going through it.

Got the modem and internet works. Need to integrate into CERI central processing.

MPHB. Withers needs to determine current orientation so we can pull it and orient correctly. Helicorders look strange. Need to verify the data.

VBMS. Vault is done. Contractor work is complete. Alena says sensors might be available in a month but NEIC staff is currently tied up with carribean net (tough job but someone has to do it).

Langston Zack. Greg's list said to test accelerometers and mass produce hardware. This is wayback burner. Mitch will check with Chuck to verify that its even going to be done anymore. Mitch asked and its unknown. Strike from list.